CH/Agen/95/03

Agenda

Charissa Group Meeting

To be held on: Friday, 24 Nov 1995, 10.30 am To be held at: Daresbury,

NSF Conference Room

Distribution: charissa@nnga.dl.ac.uk, DDW, all by e-mail/server

10.30 Introduction  BRF

Morning session — Heavy systems

10.45 ^{16}O+^{40}Ca  SMS
11.00 ^{32}S(^{16}O+^{16}O) and ^{28}Si(^{12}C+^{16}O)  NC
11.15 ^{28}Si oblate states  NC
11.30 ^{12}C on ^{24}Mg  SMS
11.45 ^{12}C(^{16}O,^{12}C^{12}C)^{4}He and ^{16}O(^{16}O,^{20}Ne)  MF
12.00 ^{12}C(^{24}Mg,^{12}C,^{12}C*)^{12}C  SH
12.15 ^{16}O+^{12}C to α + α+^{20}Ne  GRK
12.30 ^{16}O target run  RC
12.45 MEGHA module test results and Eurisys detectors RC

Afternoon session — Multiple alpha systems

14.00 Angular correlation analysis  SPGC
14.30 CC analysis of 32.5 MeV ^{12}C+^{12}C resonance  PF
15.15 ^{20}Ne+^{12}C  SH
15.30 ^{12}C+^{12}C fine structure EF  RLM
15.45 ^{12}C+^{12}C O+Be channels  PL
14.45 ^{12}C+^{12}C and ^{12}C+^{16}O excitation functions  SF/CDJ/SPGC

W.N. Catford (Secretary) 1 Nov 1995